# Visual collective behaviors on spherical robots

### v2

@@ 2.5 Robots-in-the-loop computer architecture
We propose a robots-in-the-loop flocking that consists of a set of $N$ Sphero® Bolt robots, which are devoid of visual sensor and are controlled via Bluetooth®. Their positions and headings are calculated by processing the images from an industrial 1920x1200 RGB, 164 FPS overhead camera. From all their positions, a binary visual map is recreated. This visual map serves as the input for the collective motion model. This process is shown in Fig. 6.
Figure 6: General Block diagram. The N flocking robots are seen by the 1920x1200px, 164fps camera. This image is acquired and processed to find the XY-coordinates, heading and colors for all $N$ robots. With these coordinates each robot can be identified. The XY-coordinates of the Anchor can be set at any point of the image to be integrated to the identification of the robots. All these coordinates are processed to reconstruct the binary panorama for each of the spherical robots. Each binary Panorama is the input of independent Collective motion model that process the visual panoramic information. All these models are ran in parallel and their output in term of heading are sent to each robot via a Bluetooth interface.
There are two main Python programs to handle the implementation of these blocks. The first one (administrator.py) manages the camera communication, data extraction, image processing, object identification, and binary panorama reconstruction (Fig. 3). The output is transferred to $N$ independent routines named PyRobot.py (the second program), which take a single binary panorama, perform the collective motion model, manage the Bluetooth connection, and send the desired heading and velocity (Fig. 5). Each robot receives the inputs (heading and velocity) and moves accordingly. Each PyRobot.py is a routine that can be executed either locally or remotely from the Administrator. This setup allows for scaling the flock size to the maximum number that can fit within the camera’s recognition area. Note that this scaling is hardware-constrained by the Bluetooth interface capabilities of the computer. For Windows®machines, the theoretical maximum is seven devices per machine; in Linux, it is seven per interface. However, it is recommended to have a maximum of five on the same machine managing the camera and six per interface on independent machines.
@@ 2.6 Metrics
$P(t)$ $=$ $\left\|\overline{\mathbf{e}_{i}(t)}\right\|,$ (1b)
$M(t)$ $=$ $\left\|\overline{\mathbf{y}_{i}(t)\times\mathbf{e}_{i}(t)}\right\|,$ (1c)
@@
where the overbar denotes an average over all individuals, and the unit vector $\mathbf{y}_{i}(t)=\frac{\mathbf{x}_{i}(t)-\overline{\mathbf{x}_{i}}(t)}{\|\mathbf{x}_{i}(t)-\overline{\mathbf{x}_{i}(t)}\|}$ points towards particle $i$ from the center of mass. All three metrics range within the interval $[0,1]$. The polarization $P(t)$ measures alignment: $P(t)=0$ corresponds to particles pointing in all directions, while $P(t)=1$ represents a perfectly aligned school. The milling $M(t)$ quantifies the normalized angular momentum: $M(t)=0$ corresponds to a straight-line formation, and $M(t)=1$ represents perfect milling. The opacity $O(t)$ measures the “occupancy” of the visual fields: $O(t)=0$ indicates no object in the visual field, and $O(t)=1$ signifies that the entire visual field is obscured.
@@ 3.1 Simulation of supplemented visual model
To make the problem dimensionless, we chose $a=1$ and $U=1$. We also kept the blind angle parameter and noise strength with values that lead to the most phase reproduction: $\epsilon=1$, $k_{\eta}=0.01$. With this approach, three dimensionless parameters remain: the strengths of anchor attraction and alignment, $k_{c}$, $k_{\odot}$, and $k_{\parallel}$. The anchor is a parameter aiming for a contained robotic experimentation, set practically to the minimum value that ensures the robots do not reach the arena’s border when no other law is in effect, $k_{c}=0.005$.
The visual original model has been supplemented with three major changes impacting the phase distribution and the metric on the parameter space $(k_{\odot},k_{\parallel})$:
  • the inclusion of the avoidance rule ($\mathcal{A}(\phi)$) (Fig. S7),
  • the addition of an anchor law ($\omega_{c}$) (Fig. S8),
  • the introduction of a time delay ($\tau$) (only in simulation) to reflect the time delay between the desired heading and a perceived change in heading while operating the Spheros robot, (Fig. 7).
@@
We now compare the original model with each consecutive addition by exploring the parameter space $(k_{\odot},k_{\parallel})\in[0,0.3]\times[0,0.3]$ on a flock of 10 individuals. For each parameter set, we ran 10 simulations over a long duration ($\Delta t=5000$). The mean values of $P$, $M$, and $O$ were determined by averaging over the last 1000 time units to ensure that transients do not influence the results. The outcomes of these simulations are synthesized in phase diagrams (Fig. 8).
We maintain the identification of the collective phases from the values of the polarization $P$ and milling $M$ parameters: schooling when $P>0.5$ and $M<0.5$, milling when $P<0.5$ and $M>0.5$, and swarming when $P<0.5$ and $M<0.5$.
Figure 8: Simulation Phase Diagrams for 10 individuals. The dotted line is $P=0.5$, the dashed line is $M=0.5$, and the solid line is $O=0.3$. A) Color code, b) Original visual model [28], c) Supplemented visual model with avoidance term, $\mathcal{A}(\mathcal{R}(\phi,t))$, d) Supplemented model with avoidance and anchor term, $\omega_{c}(t)$, e) Supplemented visual model with avoidance, anchoring and perceived time delay ($\tau$=2.4[a/U]).
@@ 3.2 Phase diagrams of the supplemented model in simulation
We now compare the model with each consecutive addition and the robotic implementation, exploring the parameter space $(k_{\odot},k_{\parallel})\in[0,0.3]\times[0,0.3]$ for a flock of 10 independently controlled robots. For the robotic implementation, the dimensionless parameters are set to match the radius in pixels and velocity in radius per second, resulting in $a=1$, with the other parameters remaining the same ($\epsilon=1$, $k_{\eta}=0$, and $k_{c}=0.005$). For each parameter set, we ran 5 robotic experiments over a long duration (240sec or $\Delta t\approx 8500$). The mean values of $P$, $M$, and $O$ were determined by averaging over the last 1000 time units to ensure that transients do not influence the results. The outcomes of these experiments are synthesized in phase diagrams (Fig. 9a). An example of swarming is shown in Fig. 9b, of partial schooling for a bistable phase in Fig. 9c, and of milling in Fig. 9d.
The phase diagram of the original model [28] showcases the reproduction of the main three phases (swarming, schooling, and milling), with the area above $k_{\odot}>0.15$ and $k_{\parallel}>0.15$ being bistable (Fig. 8a). A bistable phase oscillates between milling and schooling (Fig. 8b). By adding repulsion to the model, this bistable phase is removed, but the area where milling appears shrinks to roughly half of the original while the final opacity of the flocks is reduced. Opacity is an inverse metric of flock density, and this reduction aligns with the expected effect of adding avoidance.
@@ 3.3 Comparison with the phase diagram of the Spheros’ robots-in-the-loop
For the robotic implementation (see Fig. 9), we found that the milling phase remains fairly stable, appearing for high alignment ($k_{\parallel}>0.15$) across most attraction values ($k_{\odot}<0.16$). The swarming phase is constant for low alignment ($k_{\parallel}<0.06$) or high attraction ($k_{\odot}>0.17$).
@@ 4 Discussion
The collective motion model with the robotic platform reproduces several collective patterns such as swarming, milling, and a bistable phase. For the first time to the best of our knowledge, a visual flocking on spherical robots shows a phase diagram that matches almost perfectly that of the simulation with the same visual model (Fig. 8e vs Fig. 9a). The mean squared error (MSE) between the phase diagram images of robots-in-the-loop and supplemented visual model results is 2.48% (for more details see Fig. S2). Such implementation therefore bridges the gap between collective model simulation and real-world experiments, allowing for expansion to $N$ individuals with low complexity. A model based on visual information enables a mixed-reality approach where there can be interaction between the flock and the environment without adding complexity to the model or sensors available to each robot.
Future works with these spherical robots could include an explicit channel for avoidance also using optic flow, a moving anchor, a replacement of the anchor by another particle without adding a new channel, and an expansion of the flock size to 20 or 40 individuals for a closer-to-nature implementation of these visual collective models.
@@ Acknowledgements
Data acquired during the experiments and simulations are available online [41, 42]
@@
D.C. performed the simulations, the robotic experiments, and data consolidation. C.E. & F.R. outlined the robot in the loop experiment and supervised its implementation. All authors: Proofread the manuscript.
See pages 1- of SuppData2.pdf

### v3

@@ 2.5 Robots-in-the-loop computer architecture
We propose a robots-in-the-loop flocking framework that consists of a set of $N$ Sphero® Bolt robots, which lack visual sensors and are controlled via Bluetooth®. Their positions and headings are determined by processing images captured by an industrial 1920x1200 RGB, 164 FPS overhead camera. From these positions, a binary panorama is emulated. This binary panorama serves as the input to the collective motion model, as shown in Fig. 6.
Figure 6: General Block Diagram. The $N$ flocking robots are observed using a 1920x1200 px, 164 fps camera. The captured images are processed to determine the XY-coordinates, heading, and colors of all $N$ robots. Based on these coordinates, each robot is uniquely identified. The XY-coordinates of the Anchor can be set at any point within the image to assist in robot identification. These coordinates are further processed to reconstruct the binary panorama for each spherical robot. Each binary panorama serves as input to an independent collective motion model, which processes the visual panoramic data. These models operate in parallel, and their outputs, in terms of heading, are transmitted to each robot via a Bluetooth interface.
There are two main Python programs to handle the implementation of the blocks, as shown in Fig. 6. The first one (administrator.py) coordinates the camera communication, data extraction, image processing, object identification, and binary panorama reconstruction (Fig. 3). The binary panorama reconstruction is transferred to $N$ independent routines named PyRobot.py (the second program). Each routine processes the binary panorama of a single robot to calculate its output from the collective motion model and send the desired heading and speed via Bluetooth to that same robot (Fig. 5). Each robot receives the inputs (heading and velocity) and moves accordingly. Each PyRobot.py is a routine that can be executed either locally or remotely from the administrator. This setup allows a framework only bounded in the flock size to the maximum number that can fit within the camera’s recognition area and hardware-constrained Bluetooth interface capabilities of the computer. For Windows® machines, the theoretical maximum is seven devices per machine; in Linux, it is seven per interface. However, in practice, it is recommended to have a maximum of five on the same machine managing the camera and six per interface on independent machines.
@@ 2.6 Metrics
$P(t)$ $=$ $\left\|\overline{\mathbf{e}_{i}(t)}\right\|,$ (2c)
$M(t)$ $=$ $\left\|\overline{\mathbf{y}_{i}(t)\times\mathbf{e}_{i}(t)}\right\|,$ (2d)
@@
where the overbar denotes an average over all individuals, and the unit vector $\mathbf{y}_{i}(t)=\frac{\mathbf{x}_{i}(t)-\overline{\mathbf{x}_{i}}(t)}{\|\mathbf{x}_{i}(t)-\overline{\mathbf{x}_{i}(t)}\|}$ points towards particle $i$ from the center of mass. All three metrics range within the interval $[0,1]$. The polarization $P(t)$ measures alignment: $P(t)=0$ corresponds to particles pointing in all directions, while $P(t)=1$ represents a perfectly aligned school. The milling $M(t)$ quantifies the normalized angular momentum: $M(t)=0$ corresponds to a straight-line formation, and $M(t)=1$ represents perfect milling. The opacity $O(t)$ measures the “occupancy” of the visual fields: $O(t)=0$ indicates no object in the visual field, and $O(t)=1$ signifies that the entire visual field is obscured. To asses the performance of the phase, we quantify the number of collisions while a simulation or experiment were running. For an experiment, we counted how many frames presented a collision and how many robots were involved. For a simulation we counted collision by defining one as any time the distance between individuals was less than $2a$.
@@ 3.1 Simulation of Supplemented Visual Model
To make the problem dimensionless, we chose $a=1$ and $U=1$. The blind angle parameter and noise strength were set to values that enable the reproduction of most phases: $\epsilon=1$, $k_{\eta}=0.01$. With this approach, three dimensionless parameters remain: the strengths of anchor attraction and alignment, $k_{c}$, $k_{\odot}$, and $k_{\parallel}$. The anchor parameter was chosen to confine robotic experimentation, set practically to the minimum value required to prevent the robots from reaching the arena’s border when no other law is in effect, $k_{c}=0.005$.
The original visual model has been supplemented with three significant modifications, influencing the phase distribution and metrics in the parameter space $(k_{\odot},k_{\parallel})$:
  • inclusion of the avoidance rule ($\mathcal{A}(\phi)$) (Fig. S7),
  • addition of an anchor law ($\omega_{c}$) (Fig.S8),
  • introduction of a time delay ($\tau$) (simulation only) to reflect the delay between the desired heading and a perceived change in heading when operating the Spheros robot (Fig.7).
@@
We compared the original model with each successive addition by exploring the parameter space $(k_{\odot},k_{\parallel})\in[0,0.3]\times[0,0.3]$ for a flock of 10 individuals. For each parameter set, we ran 10 simulations over an extended duration ($\Delta t=5000$). The mean values of $P$, $M$, and $O$ were determined by averaging over the final 1000 time units to minimize the influence of transients. The simulation results are summarized in phase diagrams (Fig. 8).
We classified the collective phases using the polarization $P$ and milling $M$ parameters:
Schooling when $P>0.5$ and $M<0.5$, Milling when $P<0.5$ and $M>0.5$, Swarming when $P<0.5$ and $M<0.5$.
Figure 8: Simulation Phase Diagrams for 10 individuals. The dotted line indicates $P=0.5$, the dashed line indicates $M=0.5$, and the solid line indicates $O=0.3$. a) Color code, b) Original visual model [33], c) Supplemented visual model with avoidance term, $\mathcal{A}(\mathcal{R}(\phi,t))$, d) Supplemented model with avoidance and anchor term, $\omega_{c}(t)$, e) Supplemented visual model with avoidance, anchoring, and perceived time delay ($\tau=2.4[a/U]$).
@@ 3.2 Phase Diagrams of the Supplemented Model in Simulation
We further compared the supplemented model and its robotic implementation by exploring the parameter space $(k_{\odot},k_{\parallel})\in[0,0.3]\times[0,0.3]$ for a flock of 10 independently controlled robots. For robotic experiments, dimensionless parameters were adjusted to match the radius in pixels and velocity in radius per second, resulting in $a=1$, with other parameters remaining constant ($\epsilon=1$, $k_{\eta}=0$, and $k_{c}=0.005$).
For each parameter set, five robotic experiments were conducted over an extended period (240 sec or $\Delta t\approx 8500$). The mean values of $P$, $M$, and $O$ were calculated by averaging over the final 1000 time units to eliminate transient effects. The experimental results are summarized in phase diagrams (Fig.9a). Examples of swarming, partial schooling in a bistable state, and milling are presented in Fig.9b, Fig.9c, and Fig.9d, respectively. Note that the bistable state has been observed in fish experiments ([46]).
The phase diagram of the original model [33] showcases the reproduction of the main three phases (swarming, schooling, and milling), with the area above $k_{\odot}>0.15$ and $k_{\parallel}>0.15$ being bistable (Fig. 8a). A bistable state is when the flock oscillates between milling and schooling (Fig. 8b). By adding repulsion to the model, this bistable phase is removed, but the area where milling appears shrinks to roughly half of the original while the final opacity of the flocks is reduced. Opacity is an inverse metric of flock density, and this reduction aligns with the expected effect of adding avoidance.
@@ 3.3 Comparison with the phase diagram of the Spheros’ robots-in-the-loop
During the robotic implementation (see Fig. 9), the milling phase remains fairly stable, appearing for high alignment ($k_{\parallel}>0.15$) across most attraction values ($k_{\odot}<0.16$). The swarming phase is constant for low alignment ($k_{\parallel}<0.06$) or high attraction ($k_{\odot}>0.17$).
@@ 4 Discussion
In this article, we presented a robot-in-the-loop approach to a visual collective motion model. The robots used are practical to deploy and commercially available: the Sphero BOLT (Fig. 2). The Sphero BOLTs are sphere-like robots with a transparent shell that encloses all electronics, sensors, motors, batteries, and LEDs in a sealed design. However, due to this sealed design, integrating vision required an external solution. We opted for an external infrastructure made of overhead camera system to recognize the robots, enabling the reconstruction of each robot’s visual field. Each robot was assigned a unique light pattern, visible through the transparent shell, to facilitate position and heading recognition via the overhead camera. (Fig. 3).
The noise effect
The collective motion model presented by Castro et al. (2024) [33] and the supplemented version introduced in this article, incorporated noise as a component during the computation of angular velocity model output which is directly used to actuate the robot. Introducing the noise on the actuation signal accounted for the noise that errors on the heading recognition might introduce, but did not accounted for recognition errors of the position. And, by extension, the noise that those errors might introduce on the robots’ panorama reconstruction hence on the measurement of the “early vision” cues (apparent distance, optical position, and optic flow signals on Fig. 1). The effect of noise on these “early vision” cues can be particularly significant for the alignment term in the model. Alignment is calculated using optic flow and apparent distance (Eq. 2d), making it three times more sensitive to measurement noise than attraction that only uses apparent distance (Eq. 2c).
Fig. S1 illustrates the combined effect of these recognition errors in open loop, showing that the mean error is approximately 6%. Consequently, the attraction term inherently has a 6% error, while alignment error reaches around 15%. Despite this, the simulation and experimental robotic results align on a macro level as shown in Fig. 9. For comparison, the mean squared error (MSE) between the phase diagram images from robots-in-the-loop and supplemented visual model results is as low as 2.48% (for more details, see Fig. S2).
Effect of collisions
We analyzed the number of collisions throughout the evolution of the flock during the examples of swarming, bistability, and milling. If a collision occurred during the robotic experiment or simulation, we counted how many individuals participated on the collision (as seen on Fig. S9). In simulation, a collision is detected when the distance between individuals is less than or equal to $2a$. In the robotic experiment, a collision is detected when the recognized object becomes much larger than a single Sphero. These collisions are calculated throughout the simulation and the robotic experiment, and are presented as additional information on the state of the flock (Fig. S10).
We found that, on average, in one video for each phase, the flock in simulation spent 14% of their time experiencing collisions. The robots exhibited a collision rate of 33% of the time (on average, three times more than the simulation). Interestingly, most simulation collisions were between two particles, and instances of collisions involving more than two particles were rare (only seen on Fig S9-b). In contrast, with the robots, most collisions caused a chain reaction, leading other particles to collide with the stopped particles (as seen on Fig S9-d). The most collisions occurred during the bistable state, where the simulation had a 21% collision rate compared to 36% for the robots. In both cases, the fewest collisions occurred during the milling phase. The simulation only showed a 6% collision rate, while the robots exhibited a 30% collision rate. The stability of the milling phase is also seen on the minimum inter-distance that, on simulation and on the robots, tends to stabilize around t=1200. The timeline of this phase can be seen on Fig S9-c for simulation and Fig S9-f for robots.
The anchor’s effect
One notable modification to the model was the introduction of an attractive force towards the center of the area of interest, referred to as the anchor (Eq. 1). As shown in Fig. 8c, the main phase is schooling. However, when the anchor is introduced (Fig. 8d), schooling disappears. We hypothesize that the anchor, acting as a force that causes individuals to turn towards the center when they are far from it, challenges the formation of schooling. Notably, schooling is typically observed during flock translation, such as animal migration, where the flock’s center of mass moves along the flock’s polarized direction.
If the model bounds a flock with a wall, a schooling flock would inevitably interact with the wall. In contrast, our model introduces a force directing individuals toward the a desired location, the center of the flock. This means a schooling flock, while temporarily moving away, would eventually return toward the selected point. Moreover, the anchor acts as an unbiased attraction force, causing individuals to turn either clockwise or counterclockwise. During this turning, the flock depolarizes, leading to phase changes. As a result, any observed schooling is temporary.
The delay’s effect
Another significant factor contributing to the expansion of milling is the delay. A delay between recognition and actuation has been documented to influence collective motion. [47, 48, 49, 50]. If the anchor interacts with the flock with an attractive force, the delays does it with an alignment force. Because the delay, the particles tend to align with a past velocity. In other words, the polarization perceived by the particles is delayed, which leads to longer convergence on the 2 coordinated phases, only schooling or milling. This longer convergence time that the coordinated phase takes to stabilize creates opportunities for other phases to replace the current coordinated phase. Particularly the swarming phase expands because the flock itself maintains a uncoordinated cohesion on high attraction or low alignment (Fig. 8d). For the case of the bistability region, it expands on the borders of the coordinated phases, as the delay causes the flock to oscillate between milling and schooling phases more often.
Comparison between phase diagrams
The comparison between the robotic (Fig. 9a) and the simulated phase diagrams (Fig. 8d) only differs for the small area of polarized bistability. The low-attraction region differs because collisions are more relevant in the experimental case than in simulation, where avoidance was included only in the attraction mechanism (low attraction equals low avoidance). In the high-attraction region, the small difference is attributable to the significant noise present in the alignment rule, caused by the compounding effect of optic flow calculations based on retinal location and apparent distance. All three primary effects of the robotic implementation expands the area of a stable milling phase.
5 Conclusion
@@
Regardless of these implementation challenges, the collective motion model applied on the robotic platform reproduces several collective patterns, such as swarming, milling, and a bistable phase. To the best of our knowledge, for the first time, a visual flocking implementation using spherical robots produces a phase diagram that matches almost perfectly with that of the simulation counterpart using the same visual model. Such an implementation bridges the gap between collective model simulation and real-world experiments, allowing for a framework independent of the $N$ individuals belonging to the flock.
Future Work
For future work, it would be interesting to implement an obstacle avoidance rule based on optic flow. With the divergence of the optic flow, potential collisions could be better avoided. It would also be valuable to explore how the presence of the most influential retinal object could affect the phase distribution or evaluate the current model to identify the average location of the most influential retinal object. Another interesting approach would be to introduce a true blind spot where there is no sensing. Potential applications include the evaluation of a moving anchor or transforming the concept of the anchor into a “guarding dog” individual—one individual that moves around and keeps the flock in the desired position—or to include the anchor as a static individual. Another interesting path would be the implementation to 20 or 40 individuals, or even incorporating a mixed-reality option with virtual individuals that interact with the robots to create flocks of hundreds.
@@ Acknowledgements
We would also like to thank the two reviewers for their valuable suggestions and comments.
@@
Data acquired during the experiments and simulations are available online [51]
@@
D.C. performed the simulations, the robotic experiments, and data consolidation. C.E. & F.R. outlined the robot in the loop experiment and supervised its implementation. All authors: Proofread the manuscript. See pages 1- of SuppData2.pdf